# A Context-Driven Approach for Co-Auditing Smart Contracts with The Support of GPT-4 code interpreter

### v1

@@ III-A2 Function Call Graph (FCG) Extraction
The single file resulting from the flattening process is long and cannot fit within one prompt. One way to provide GPT with enough context is to focus on the Function Call Graph (FCG), which shows the flow of function calls, indicating which functions are called by other functions and how they are connected. FCG is utilized in our co-auditing approach to aid the segmentation mechanism to send only logically related code segments based on inter-dependencies rather than their physical location within the codebase. In our work, to generate a FCG from the flattened solidity file, we again used Slither to extract function calls from each contract. Slither uses the context of function calls and applies static and data flow analysis techniques to distinguish between internal and external function calls. The extracted function calls are then consolidated into a call graph.
III-A3 CCL Generation
@@
Based on the FCG, we generate the CCL for each function. A CCL is a list containing all the function codes that are called during a specific function execution, which provides the GPT-4 code interpreter with a self-containing list of function codes. The number of CCLs corresponds to the number of functions implemented within the smart contract. Each CCL is composed of the code of the audited function, the caller function, and the code of the callee functions. This process is also recursive for the callee functions. This process guarantees that all the code being executed during a specific function call is included within the prompt. Our approach differentiates between external and internal function calls in the contracts. Using the function name and the type of the call, we locate and extract the exact called function and its corresponding code. This aims to provide the LLM with enough context to fully process the functions and recognize the vulnerability patterns.
III-B Reporting Scoping - CAQ Prompt Design
@@
The effectiveness of LLMs in audit reporting heavily relies on how specifically the questions are posed. Accurate and relevant responses from LLMs depend on question precision, which is crucial in audit reporting, where delivering exact content in a certain order and format is essential. By formulating targeted questions, we direct the LLM to produce the required content accurately and focus on generating subsequent relevant information. Therefore, integrating a response template into the prompt is a key strategy to reduce response variability and standardize the output format of LLMs in audit reporting. In order to do this, we adopted a focus group methodology to identify the essential components of audit reports. Initially, a diverse collection of existing audit reports from our partner and from Code4rena [^2] https://code4rena.com/reports is assembled for analysis to understand the structure and content of audit reports. Then, a focused group of two professional auditors, in addition to three students, examined the reports. Each member of the group was asked to extract a schema of the main report parts. The group then agreed on a common schema that represents key report components that need to be included in the audit report. Based on the discussion, the group agreed that a report needs to specify (i) if a vulnerability exists, (ii) a proof of the concept of the vulnerability, (iii) the business impact of the vulnerability, and (iv) suggestions for potential fixes. Building on this, we create the CAQ-based prompt strategy, then refine it to address Assessment Scoping, as we explain in the next section. The CAQ-based prompt strategy aims at eliciting specific, relevant responses in alignment with these identified components. An initial assessment of the proposed prompt template. Pilot testing and iterative refinement based on feedback from both the professional auditors and the students were conducted on a small dataset, leading to a finalized template that generated acceptable responses. To limit the impact of the non-deterministic nature of generative models, we fixed the temperature in all tests to 0 in the GPT-code-interpreter during the pilot testing. Figure  2 illustrates the CAQ-based prompt structure. To get a complete audit report based on this approach, the number of prompt executions is equal to the number of functions in a smart contract.
@@
III-C Assessment Scoping - CWE Prompt Design
@@
Figure  3 illustrates the structure for CWE-based prompts. For each function, we execute this prompt for all 38 vulnerabilities identified in the CWE. In order to generate a complete audit report for a smart contract, we repeat the execution of the prompt for every function in the contract. The total prompt execution required equals 38 multiplied by the number of functions in the contract. This prompt design ensures we thoroughly evaluate the contract for known vulnerabilities and provide a comprehensive audit report.
@@ IV-A Data Sources
The vulnerable smart contracts in both datasets are classified into different categories. Each smart contract belongs to only one category, and only one function of the smart contract is vulnerable. During our preliminary study, we noticed that GPT reveals different vulnerabilities in one smart contract. These vulnerabilities affect functions that are not claimed to be vulnerable. Moreover, although these datasets contain the labeled vulnerabilities, they might also include other vulnerable functions that are not labeled. To ensure the integrity of the study and given that GPT could have seen the publically available labeled datasets, we considered unlabeled functions in datasets. Hence, in this study, we divide the data sources into two datasets: Dataset A and Dataset B. Dataset A includes the labeled vulnerable functions, while Dataset B has unlabeled functions extracted from the data sources. The dataset B is composed only of unlabeled functions.
IV-B Experiment Setup
TABLE I: Experiment Sources
@@ IV-B Experiment Setup
As mentioned in section III, our approach is based on three scoping strategies: code scoping aims to shorten the prompted code, reporting scoping, provides GPT with a standardized output format for the audits, and assessment scoping, sets a focus at the targeted vulnerability that GPT should address. These strategies were used to guide the design of different prompting templates. In our empirical study, we study the impact of employing these strategies on the GPT performance in smart contract code auditing.
We designed six experiments to answer the research questions based on the different prompting templates and whether they are used with the full code or the Code Call List (CLL). Table I shows all the experiments; columns show the source name as it appears in the replication repository, the dataset used, the prompt template used, and whether the experiment is based on the full code of CCL, respectively. These experiments were used to answer the questions listed in section II
TABLE II: Dataset A experimental results
@@
IV-C Empirical Study Results
IV-C1 Impact of the code scoping on the GPT performance
@@
After prompting GPT-4 interpreter using the different strategies, the responses are collected in CSV files and manually compared to the ground truth. A response is considered correct if GPT correctly identifies the same type and exact location of the vulnerability as described in the ground truth contracts. For each type of vulnerability, we report the number of correctly detected vulnerabilities as shown in table II.
To answer RQ1.1, we conduct hypothesis testing[18] on results from sources 1 and 3 in Table  4. We formulated the null hypothesis (H0) and the alternative hypothesis (H1) as follows:
@@
To compare the efficacy of these strategies (i.e., CAQ prompt with CCL-based chunking vs. CAQ prompt with full code), which operate as classifiers, we use McNemar’s Test, which is particularly useful for comparing the performance of two classifiers. The test statics measurement indicates a value of 62.01 following a chi-squared distribution with one degree of freedom. For a significance level of 0.05, consulting the chi-squared distribution table for one degree of freedom, we identified a critical value of 3.841. Hence, the test statics are greater than the critical value; we can conclude that there is a significant difference in terms of effectiveness between the full code and CCL-based chunking methods.
The detection rates are 59.24% and 94.02% for full code and CCL-based chunking methods, respectively. The cohen’s d value is equal to -0.45 which suggests that the difference in effectiveness between the two approaches is moderately significant. The stark contrast in these central tendencies, coupled with a notably low p-value of approximately 0.000287, strongly supports the rejection of the RQ1.1-H0. Therefore, we conclude that when utilizing the CAQ prompt, there are differences in the performance of CCL-based chunking compared to that of full code prompting, with CCL-based chunking performing much better.
To answer RQ1.2, we conducted a hypothesis test on the results from sources 2 and 4 as presented in Table  I. We formulated the null hypothesis (H0) and the alternative hypothesis (H1) as follows:
@@
Similar to RQ1.1, we compare the efficacy of these strategies (i.e., CWE prompt with CCL-based chunking vs. CWE prompt with full code) by using McNemar’s Test. Applying the test static formula of McNemar’s Test indicates a value of 50.01, which is superior to the critical value obtained from the chi-squared distribution table with one degree of freedom, which is equal to 3.841. This value indicates a significant performance difference between the two approaches. The Cohen’s d-factor is measured at -0.36, suggesting that the mean of correctly detected vulnerabilities is lower for full code prompting compared to CCL-based chunking. The substantial disparity between these means is robustly supported by a p-value of approximately 0.00117, surpassing conventional thresholds for statistical significance and compellingly justifying the rejection of the null hypothesis. The detection rate is 67.93% and 96.02% respectively. Also the sensitivity, F1-score and precision are 76.69%, 80.91%, 85.62% and 98.88%, 98.06%, 97.25% respectively. Consequently, we assert that under the CWE prompt, differences exist in the performance of CCL-based chunking and full code prompting, with CCL-based chunking demonstrating superior performance.
@@
IV-C2 Impact of the assessment scoping on GPT performance
@@
In order to assess and compare the performance of the CAQ and CWE prompts, we calculate the McNemar’s Test [19], which results in a value equal to 2.25. This value is inferior to the critical value obtained from the chi-squared distribution table for 1 degree of freedom and 0.05 confidence significance level. Using this metric, we are not able to reject the Null Hypothesis, RQ2-H0. The statistical analysis yielded a p-value of approximately 0.553, which fails to provide substantial evidence to reject the null hypothesis, indicating no significant difference. To calculate the effect size, we first calculate the standard deviation for CAQ and CWE prompts, which result in 15.61 and 16.04, respectively. The Cohen’s d value resulted is -0.025, which is considered very tight. This value indicates that there is a very small difference between the two approaches in terms of their ability to detect vulnerabilities.
@@
IV-C3 Human qualitative evaluations
@@
Addressing RQ3, we aim to evaluate the reliability and add an extra layer of validity to the proposed approach. We conducted a qualitative experiment in which we asked expert human auditors to audit and manually validate the auditing reports generated by our approach on a set of unlabeled smart contract functions. This evaluation was conducted under the assumption of a low likelihood that GPT models being previously exposed to, or trained on, the specific function annotations (labels). We hope this will provide insights into how well our approach performs in scenarios likely unfamiliar to the GPT’s training data. The following sections detail the dataset, the experiment and the insights from validation.
@@
We employed a rigorous annotation process for the manual study of GPT-generated audits of the selected 86 CCL. Two experienced auditors from our industrial partner annotated the samples independently along the two dimensions. The annotators have extensive expertise in the development and auditing of smart contract projects. The process of the annotation of the GPT responses is held into rounds. First, each auditor independently reviewed and evaluated the GPT response based on the correctness and relevance of the response. Then, a meeting between the two annotators and the co-authors is held to ensure they share the same understanding of the task. The goal is to enable the annotators to perform their tasks with a well-defined framework and to enhance the quality and consistency of their individual assessments, recognizing that differences in judgment are inherent in this challenging domain.
The annotators were able to access the full source code corresponding to each CCL in order to have an accurate understanding of the functions being audited. Different meetings were held between the auditors in order to solve the discrepancies between annotations. The final Cohen’s Kappa coefficient [20] for response correctness and response relevance was 0.65 and 0.68, respectively, suggesting moderate agreement between the two annotators. Even though the relevance score between the two annotators is not exceptionally high, the labeling results retain their value. This is because our analysis focuses on comparing the changes in scores assigned by the same annotator to the CCL-based chunking and full code usage performances rather than making a direct comparison between the scores from two annotators. This approach underscores the individual annotator’s consistency and personal benchmark standards, providing a more controlled and internally consistent comparison.
@@
Results: The outcomes of the annotation procedure are depicted in Figure 5. The results show despite using unlabeled datasets, on average, expert auditors concurred that with our prompting approach, GPT-4 interpreter correctly and accurately generated reports in 56% of the cases, and these reports were relevant in 46% of instances. These percentages increase to 76.5% and 75% for partial correctness and relevance, respectively. This aligns with our prior findings, that the application of CCL-based chunking enhances the GPT-4 code interpreter’s ability to identify vulnerabilities in unlabeled datasets.The derived auditing dataset can be accessed openly in the project’s repository.  [^4] https://zenodo.org/records/10530868

### v2

@@ 3.1.2 Function Call Graph (FCG) Extraction
The single file resulting from the flattening process is long and cannot fit within one prompt. One way to provide GPT with enough context is to focus on the Function Call Graph (FCG), which shows the flow of function calls, indicating which functions are called by other functions and how they are connected. FCG is utilized in our co-auditing approach to aid the segmentation mechanism to send only logically related code segments based on inter-dependencies rather than their physical location within the codebase. In our work, to generate a FCG from the flattened solidity file, we again used Slither to extract function calls from each contract. Slither uses the context of function calls and applies static and data flow analysis techniques to distinguish between internal and external function calls. The extracted function calls are then consolidated into a call graph. Based on the smart contract code in Section  2.3, the generated FCG is shown in figure  2.
Figure 2: Function call graph of the tokenizeVault contract
3.1.3 CCL Generation
Algorithm 1 CCL generation
Input: flattenedContract, adjacencyList
Output: codeCallList
1: Initialize codeCallList as an empty dictionary
2:for each functionId, callRelationships in adjacencyList do
3:  for each callRelationship in callRelationships do
4:   callerFunctionId, calledFunctionId $\leftarrow$ callRelationship
5:   functionCode $\leftarrow$ extractCode(calledFunction, flattenedContract)
6:   codeCallList[functionId].append(functionCode)
7:  end for
8:end for
9:return codeCallList
@@ 3.1.3 CCL Generation
Based on the FCG, we generate the CCL for each function. A CCL is a list containing all the function codes that are called during a specific function execution, which provides the GPT-4 code interpreter with a self-containing list of function codes. The number of CCLs corresponds to the number of functions implemented within the smart contract. Each CCL is composed of the code of the audited function, the caller function, and the code of the callee functions. This process is also recursive for the callee functions. This process guarantees that all the code being executed during a specific function call is included within the prompt. Our approach differentiates between external and internal function calls in the contracts. Using the function name and the type of the call, we locate and extract the exact called function and its corresponding code. This aims to provide the LLM with enough context to fully process the functions and recognize the vulnerability patterns. The listing 2 illustrates a sample of CCL generated from the FCG for the function deploy.
Listing 2: CCL list of the function ’deploy’ extracted from contract TokenizingVault
['function deploy( ERC20 underlying ) external nonReentrant returns (ERC20CreditToken) { // Create credit token if one doesnt already exist ERC20CreditToken creditToken = creditTokens[underlying]; if (creditToken == ERC20CreditToken(address(0))) { bytes memory tokenData = abi.encodePacked( underlying, address(this) ); creditToken = ERC20CreditToken( address(creditTokenImpl).clone(tokenData) ); creditTokens[underlying] = creditToken; } return creditToken; }',
'constructor(string memory name, string memory symbol, ERC20 _underlying) ERC20(name, symbol) { underlying = _underlying; vault = msg.sender; }',
'modifier nonReentrant() { _guardCounter += 1; uint256 localCounter = _guardCounter; _; require(localCounter == _guardCounter); }']
3.2 Reporting Scoping - CAQ Prompt Design
@@
The effectiveness of LLMs in audit reporting heavily relies on how specifically the questions are posed. Accurate and relevant responses from LLMs depend on question precision, which is crucial in audit reporting, where delivering exact content in a certain order and format is essential. By formulating targeted questions, we direct the LLM to produce the required content accurately and focus on generating subsequent relevant information. Therefore, integrating a response template into the prompt is a key strategy to reduce response variability and standardize the output format of LLMs in audit reporting. In order to do this, we adopted a focus group methodology to identify the essential components of audit reports  [^2] https://code4rena.com/reports/2024-03-dittoeth. Initially, a diverse collection of existing audit reports from our partner and from Code4rena [^3] https://code4rena.com/reports is assembled for analysis to understand the structure and content of audit reports. Then, a focused group of two professional auditors, in addition to three students, examined the reports. Each member of the group was asked to extract a schema of the main report parts. The group then agreed on a common schema that represents key report components that need to be included in the audit report. Based on the discussion, the group agreed that a report needs to specify (i) if a vulnerability exists, (ii) a proof of the concept of the vulnerability, (iii) the business impact of the vulnerability, and (iv) suggestions for potential fixes. Building on this, we create the CAQ-based prompt strategy, then refine it to address Assessment Scoping, as we explain in the next section. The CAQ-based prompt strategy aims at eliciting specific, relevant responses in alignment with these identified components. Pilot testing and iterative refinement based on feedback from both the professional auditors and the students were conducted on a small dataset, leading to a finalized template that generated required, human-like responses. To limit the impact of the non-deterministic nature of generative models, we fixed the temperature in all tests to 0 in the GPT-code-interpreter during the pilot testing. Figure  3 illustrates the CAQ-based prompt structure. To get a complete audit report based on this approach, the number of prompt executions is equal to the number of functions in a smart contract.
@@
3.3 Assessment Scoping - CWE Prompt Design
@@
Figure 4 illustrates the structure for CWE-based prompts. For each function, we execute this prompt for all 38 vulnerabilities identified in the CWE. In order to generate a complete audit report for a smart contract, we repeat the execution of the prompt for every function in the contract. The total prompt execution required equals 38 multiplied by the number of functions in the contract.
@@ 4.1 Data Sources
The vulnerable smart contracts in both datasets are classified into different categories. Each smart contract belongs to only one category, and only one function of the smart contract is vulnerable. During our preliminary study, we noticed that GPT reveals different vulnerabilities in one smart contract. These vulnerabilities affect functions that are not claimed to be vulnerable. Moreover, although these datasets contain the labeled vulnerabilities, they might also include other vulnerable functions that are not labeled. To ensure the integrity of the study and given that GPT could have seen the publically available labeled datasets, we considered unlabeled functions in datasets. Hence, in this study, we divide the data sources into two datasets: Dataset A and Dataset B. Dataset A includes the labeled vulnerable functions, while Dataset B has unlabeled functions extracted from the data sources.
4.2 Experiment Setup
Table 1: Experiment Sources
@@ 4.2 Experiment Setup
As mentioned in section 3, our approach is based on three scoping strategies: code scoping aims to shorten the prompted code, reporting scoping, provides GPT with a standardized output format for the audits, and assessment scoping, sets a focus at the targeted vulnerability that GPT should address. These strategies were used to guide the design of different prompting templates. In our empirical study, we study the impact of employing these strategies on the GPT performance in smart contract code auditing.
We designed six experiments to answer the research questions based on the different prompting templates and whether they are used with the full code or the Code Call List (CLL). Table 1 shows all the experiments; columns show the source name as it appears in the replication repository, the dataset used, the prompt template used, and whether the experiment is based on the full code of CCL, respectively. These experiments were used to answer the questions listed in section 2
Table 2: Dataset A experimental results
@@
4.3 Empirical Study Results
4.3.1 Impact of the code scoping on the GPT performance
@@
After prompting GPT-4 interpreter using the different strategies, the responses are collected in CSV files and manually compared to the ground truth. A response is considered correct if GPT correctly identifies the same type and exact location of the vulnerability as described in the ground truth contracts. For each type of vulnerability, we report the number of correctly detected vulnerabilities as shown in table 2.
To answer RQ1.1, we conduct hypothesis testing[35] on results from sources 1 and 3 in Table  5. We formulated the null hypothesis (H0) and the alternative hypothesis (H1) as follows:
@@
To compare the efficiency of these strategies (i.e., CAQ prompt with CCL-based chunking vs. CAQ prompt with full code), which operate as classifiers, we use McNemar’s test, which is particularly useful for comparing the performance of two classifiers. The test statistic measurement indicates a value of 62.01 following a chi-squared distribution with one degree of freedom. For a significance level of 0.05, consulting the chi-squared distribution table for one degree of freedom, we identified a critical value of 3.841. Hence, the test statistic is greater than the critical value; we can conclude that there is a significant difference in terms of effectiveness between the full code and CCL-based chunking methods.
The detection rates are 59.24% and 94.02% for full code and CCL-based chunking methods, respectively. The Cohen’s d value is equal to -0.45, which suggests that the difference in effectiveness between the two approaches is moderately significant. The stark contrast in these central tendencies, coupled with a notably low p-value of approximately 0.000287, strongly supports the rejection of RQ1.1-H0. Therefore, we conclude that when utilizing the CAQ prompt, there are differences in the performance of CCL-based chunking compared to that of full code prompting, with CCL-based chunking performing much better.
To answer RQ1.2, we conducted a hypothesis test on the results from sources 2 and 4 as presented in Table  1. We formulated the null hypothesis (H0) and the alternative hypothesis (H1) as follows:
@@
Similar to RQ1.1, we compare the efficacy of these strategies (i.e., CWE prompt with CCL-based chunking vs. CWE prompt with full code) by using McNemar’s test. Applying the test statistic formula of McNemar’s test indicates a value of 50.01, which is superior to the critical value obtained from the chi-squared distribution table with one degree of freedom, which is equal to 3.841. This value indicates a significant performance difference between the two approaches. The Cohen’s d value is measured at -0.36, suggesting that the mean of correctly detected vulnerabilities is lower for full code prompting compared to CCL-based chunking. The substantial disparity between these means is robustly supported by a p-value of approximately 0.00117, surpassing conventional thresholds for statistical significance and compellingly justifying the rejection of the null hypothesis. The accuracy rates are 67.93% and 96.02%, respectively. Additionally, the sensitivity, F1-score, and precision are 76.69%, 80.91%, 85.62% and 98.88%, 98.06%, 97.25%, respectively. Consequently, we assert that under the CWE prompt, differences exist in the performance of CCL-based chunking and full code prompting, with CCL-based chunking demonstrating superior performance.
@@
4.3.2 Impact of the assessment scoping on GPT performance
@@
In order to assess and compare the performance of the CAQ and CWE prompts, we calculate McNemar’s test [27], which results in a value equal to 2.25. This value is inferior to the critical value obtained from the chi-squared distribution table for 1 degree of freedom and a 0.05 confidence significance level. Using this metric, we are not able to reject the null hypothesis, RQ2-H0. The statistical analysis yielded a p-value of approximately 0.553, which fails to provide substantial evidence to reject the null hypothesis, indicating no significant difference. The accuracy rates are 94.02% and 96.02%, respectively. the recall value is equal to 94.02% and 96.02%. To calculate the effect size, we first calculate the standard deviation for the CAQ and CWE prompts, which results in 15.61 and 16.04, respectively. The Cohen’s d value resulted in -0.025, which is considered very small. This value indicates that there is a very small difference between the two approaches in terms of their ability to detect vulnerabilities. Additionally, the sensitivity, F1-score, and precision are 76.69%, 80.91%, 85.62% and 98.88%, 98.06%, 97.25%,
@@
We have redone the tests on the redeem function that we explained in the section 2.3, using the CAQ-based prompt along with CCL, GPT4 was able to detect effectively the vulnerability  [^5] https://chat.openai.com/share/fb51d632-c6aa-47b0-a81e-7ec473c28b92.
4.3.3 Human qualitative evaluations
@@
Addressing RQ3, we aim to evaluate the reliability and add an extra layer of validity to the proposed approach. We conducted a qualitative experiment in which we asked expert human auditors to audit and manually validate the auditing reports generated by our approach on a set of unlabeled smart contract functions. This evaluation was conducted under the assumption of a low likelihood that GPT models were previously exposed to, or trained on, the specific function annotations (labels). We hope this will provide insights into how well our approach performs in scenarios likely unfamiliar to the GPT’s training data. The following sections detail the dataset, the experiment, and the insights from validation.
@@
Dataset B comprises unlabeled functions paired with corresponding Contract Contextual Labels (CCL). To ensure rigor, we selected contracts meeting specific criteria:
  • 1.
    Inter-contract dependencies (importing libraries/contracts): This criterion ensures the dataset captures scenarios where functions interact with external libraries or contracts, reflecting real-world complexity.
  • 2.
    Complex logic and lengthy code (more then 100 lines): This criteria is used to ensure complex contracts functionalities.
  • 3.
    Varied contract lengths (representative sampling): Including contracts of different sizes ensures the dataset covers diverse use cases, providing a balanced evaluation of model performance across both simple and complex scenarios.
  • 4.
    Developer-commented functions (simulating human auditing): Human auditors generally do audits to commented contracts. The comments help auditor to understand the code functionality and thus help him detect logic vulnerabilities.
This diversified dataset challenges Large Language Models (LLMs) and mirrors real-world auditing scenarios.
@@
We employed a rigorous annotation process for the manual study of GPT-generated audits of the selected 86 CCLs. Two experienced auditors from our industrial partner annotated the samples independently along two dimensions. The annotators have extensive expertise in the development and auditing of smart contract projects. The process of annotating the GPT responses was conducted in rounds. First, each auditor independently reviewed and evaluated the GPT responses based on the correctness and relevance of the responses. Then, a meeting between the two annotators and the co-authors was held to ensure they shared the same understanding of the task. The goal was to enable the annotators to perform their tasks with a well-defined framework and to enhance the quality and consistency of their individual assessments, recognizing that differences in judgment are inherent in this challenging domain. The annotators were able to access the full source code corresponding to each CCL to have an accurate understanding of the functions being audited. Different meetings were held between the auditors to resolve discrepancies between annotations. The final Cohen’s Kappa coefficient [22] for response correctness and response relevance was 0.65 and 0.68, respectively, suggesting moderate agreement between the two annotators. Even though the relevance score between the two annotators is not exceptionally high, the labeling results retain their value. This is because our analysis focuses on comparing the changes in scores assigned by the same annotator to the CCL-based chunking and full code usage performances rather than making a direct comparison between the scores from two annotators. This approach underscores the individual annotator’s consistency and personal benchmark standards, providing a more controlled and internally consistent comparison.
@@
Results: The outcomes of the annotation procedure are depicted in Figure 6. The results show that despite using unlabeled datasets, on average, expert auditors concurred that with our prompting approach, the GPT-4 interpreter correctly and accurately generated reports in 56% of the cases, and these reports were relevant in 46% of instances. These percentages increase to 76.5% and 75% for partial correctness and relevance, respectively. This aligns with our prior findings that the application of CCL-based chunking enhances the GPT-4 code interpreter’s ability to identify vulnerabilities in unlabeled datasets. The derived auditing dataset can be accessed openly in the project’s repository.  [^6] https://zenodo.org/records/10530868